OpenStudy (anonymous):

Mel is making a paper tepee, as shown below. What is the area of the paper required to make the lateral surface of the tepee? 37.7 cm2 69.1 cm2 603.3 cm2 301.7 cm2 it would help alot if someone explained it to me i do not get it?:/

OpenStudy (anonymous):

here is the picture

OpenStudy (slaaibak):

Do you know the formula for the surface area of a cylinder pyramid thing?

OpenStudy (anonymous):

no

OpenStudy (slaaibak):

oh mybad, you call it a cone. The formula is \[Surface Area = \pi \times radius^2 + \pi \times radius * height\]

OpenStudy (anonymous):

Rae well the formula for surface area of a cone is 1/2 [pi x D x slant height] so plug in your numbers and solve

OpenStudy (anonymous):

1 sec

OpenStudy (slaaibak):

(height is the slant height in my formula) Dakota I think you are confused. That's not the formula.

OpenStudy (anonymous):

would the radius be 12 or 6?

OpenStudy (slaaibak):

6

OpenStudy (anonymous):

well mine gets one of the answers offered yours doesnt try working it out slaaibak

OpenStudy (anonymous):

i got 414.48 that answwer isnt up there

OpenStudy (anonymous):

Rae try my method plz i got the right answer!

OpenStudy (anonymous):

ok 1 sec

OpenStudy (slaaibak):

They don't include the bottom surface it seems. Then just don't add the bottom surface. It reduces to Area = pi * radius * slant height

OpenStudy (slaaibak):

*surface area

OpenStudy (anonymous):

what is D?

OpenStudy (anonymous):

Therefore i am not confused slaaibak ;) and d is diameter

OpenStudy (anonymous):

i mean what is the diameter

OpenStudy (slaaibak):

I agree, but you stated the "surface area of a cone is" which is false. But I see your point. Lateral surface. Diamater is twice the radius.

OpenStudy (anonymous):

ok i got 113.04 can i just get hte answer

OpenStudy (anonymous):

it should be 1/2[ pi x 12 x 16} pi = 3.14

OpenStudy (anonymous):

i have tried to figure this out 50 times and i cannot get the right answer

OpenStudy (anonymous):

ok how do i get the answer from here: 1/2 (226.08)

OpenStudy (anonymous):

just multiply everything out and leave the pi until the very end then multiply by pi

OpenStudy (anonymous):

can i just get the answer now i got 1205.76

OpenStudy (anonymous):

omg [12 x 16pi] =192pi then 1/2[192] = 96pi then (96)(3.14) solve! i cant just give u the answer work it out

OpenStudy (anonymous):

u need help figure it out i just worked it out for you! geez its not rocket science just multiply!!!!!!!!!!

OpenStudy (anonymous):

and i have the answer i got it before your last thing with the equations

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

slaaibak what are you trying to write?